Find the slope -interception equation of the line that has the given characteristics : slope 0 and y-intercept (0,7)
y=
1
Expert's answer
2012-03-09T08:08:47-0500
General form of straight line equation in slope-intercept form is
y=mx+b
where m-slope, b - y-intercept
so we have
y=0*x+7=7